student measures the time T for one complete oscillation of a pendulum of length 1. Her results are shown in the table. Um TIs 0.420 + 0.001 1.3 + 0.1 She uses the formula T= to calculate the acceleration of free fall g. What is the best estimate of the percentage uncertainty in the value of g?


1
Expert's answer
2021-05-05T18:04:49-0400
"\frac{\Delta g}{g}=\frac{\Delta T}{T}+2\frac{\Delta l}{l}\\
\frac{\Delta g}{g}=\frac{0.001}{0.42}+2\frac{0.1}{1.3}=0.16=16\%"